OpenStudy (anonymous):

Find the exact value of cos 7π/12 using half-angle identities.

OpenStudy (anonymous):

it is half of \(\frac{7\pi}{6}\)

OpenStudy (anonymous):

use \[\cos(\frac{7\pi}{12})=\pm\sqrt{\frac{1+\cos(\frac{7\pi}{6})}{2}}\]

OpenStudy (anonymous):

although since you are in quadrant 2, it is negative, not positive

OpenStudy (anonymous):

so -(sqrt) 2-(sqrt(3))/2

OpenStudy (anonymous):

so do i have the right answer